OpenStudy (anonymous):

What is the energy change that will occur when 25.5 grams of oxygen gas (O2) react with excess methane (CH4) according to the reaction below? CH4 + 2 O2 CO2 + 2 H2O H = -889 kJ/mol 354 kJ of energy will be given off by the reaction 354 kJ of energy will be absorbed by the reaction 708 kJ of energy will be given off by the reaction 708 kJ of energy will be absorbed by the reaction

OpenStudy (anonymous):

ok so first you calculate what amount of O2 you have from 25,5 g: n=m/M = 25,5g / 32 gmol-1 =0,79 = 0,8 mol now you look at the raction you have written and see that in reaction there are 2 moles of O2 that react so now you calculate your share from your amount you have: 0,8/2= 0,4 = 40% so energy you need is 40 % of 889 kJ which is: 0,4*889=355,6 kJ / mol